# PDV: Prompt Directional Vectors for Zero-shot Composed Image Retrieval

### v2

@@ Abstract
Zero-shot composed image retrieval (ZS-CIR) enables image search using a reference image and text prompt without requiring specialized text-image composition networks trained on large-scale paired data. However, current ZS-CIR approaches face three critical limitations in their reliance on composed text embeddings: static query embedding representations, insufficient utilization of image embeddings, and suboptimal performance when fusing text and image embeddings. To address these challenges, we introduce the Prompt Directional Vector (PDV), a simple yet effective training-free enhancement that captures semantic modifications induced by user prompts. PDV enables three key improvements: (1) dynamic composed text embeddings where prompt adjustments are controllable via a scaling factor, (2) composed image embeddings through semantic transfer from text prompts to image features, and (3) weighted fusion of composed text and image embeddings that enhances retrieval by balancing visual and semantic similarity. Our approach serves as a plug-and-play enhancement for existing ZS-CIR methods with minimal computational overhead. Extensive experiments across multiple benchmarks demonstrate that PDV consistently improves retrieval performance when integrated with state-of-the-art ZS-CIR approaches, particularly for methods that generate accurate compositional embeddings. The code will be publicly available.
@@ 1 Introduction
Figure 1: Overview of Prompt Directional Vector (PDV) for Zero-Shot Composed Image Retrieval (ZS-CIR). (a) Standard ZS-CIR pipeline. (b) PDV calculation process. (c) Dynamic text embedding composition using PDV. (d) Fusion of composed embeddings: PDV-modified image embedding combined with composed text embedding.
@@
Gap 1: Staticly Composed Text Embedding. Target images, $I_{target}$, may not appear in the Top-K retrieved results when other gallery images produce embeddings closer to the composed text embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ than the target image embedding $\Psi_{I}(I_{target})$, as shown in Figure 1a. In such cases, users must iteratively refine their prompts and regenerate composed text embeddings, incurring additional manual effort and computational overhead.
Gap 2: Underutilisation of Reference Image Embedding. Current methods generally do not utilize the embedding $\Psi_{I}(I_{ref})$ of the reference image directly for retrieval; instead, $\Psi_{I}(I_{ref})$ is used solely for composition. This omission stems from consistently poor retrieval performance when incorporating these embeddings, as documented in multiple studies [18, 11, 2].
Gap 3: Suboptimal Performance of Image-Text Embedding Fusion. While the fusion of image and text embeddings outperforms single-modality approaches (image-only or text-only) [18, 11, 2], it still underperforms compared to composed text embeddings.
@@
PDV: Addressing Gap 3. Lastly, several studies demonstrate that the direct fusion of image and text embeddings outperforms using either input feautre (image or text-prompt) alone. [18, 11, 2]. However, this fusion approach still underperforms compared to using the composed text embeddings. This performance gap exists because prompt embeddings are significantly changed by incorporating context from the reference image. Specifically, $\Delta$ is not equivalent to $\Psi_{T}(P)$. To address this, we propose fusing the composed text and composed image embeddings, as illustrated in Figure 1d. Through varying the fusion weight factor $\beta$, we can dynamically control the balance between visual similarity to the reference image and semantic alignment with the prompt without needing to craft new prompts, or modify reference images. Lower $\beta$ values prioritize visual fidelity, while higher values emphasize semantic modifications specified in the prompt.
PDV serves as a plug-and-play enhancement for most ZS-CIR approaches, offering a simple and training-free solution. The computational overhead is minimal, requiring only the calculation of text and image embeddings from the reference image. We evaluate PDV by integrating it with four distinct ZS-CIR methods across various CIR benchmarks. Our experimental results demonstrate that all three use cases of PDV consistently improve upon the baseline approaches, particularly when the baseline method already generates accurate compositional embeddings.
Contributions. Our main contributions are as follows:
  • We introduce the Prompt Directional Vector (PDV), a simple and training-free enhancement that overcomes limitations of current Zero-Shot CIR methods.
  • We propose three novel applications of PDV: (1) dynamic composed text embedding synthesis through PDV scaling, which offers enhanced control over retrieval results without tedious prompt modification; (2) composed image embedding synthesis via semantic transfer of prompts to visual features through PDV addition, which prioritizes visual similarity; and (3) effective fusion of composed text and image embeddings, which improves overall performance and enables controllable balancing of visual and semantic similarity.
  • Through extensive experiments on multiple benchmarks with four ZS-CIR methods, we demonstrate that PDV consistently improves retrieval performance with minimal computational overhead.
@@ 3.1 Baseline ZS-CIR Framework
If $I_{target}\notin\mathbb{I}_{top-k}$, the user must reformulate the prompt and repeat the feature extraction process to obtain alternative retrieval results, incurring time and computational resource costs. Notably, as shown in Eq. 1, only the composed feature embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ directly influences the computation of $\mathbb{I}_{top-k}$ results. Although the gallery images are represented by their image embeddings, the image embedding of the reference image $\Psi_{I}(I_{ref})$ does not contribute to the retrieval process.
@@ 3.3 Strategies for Using PDV
Figure 2: Comparison of Image + Text (a) vs PDV (b).
@@
(a) PDV-T: Impact of $\alpha$ scaling on composed text embeddings
(b) PDV-I: Impact of $\alpha$ scaling on composed image embeddings
(c) PDV-F: Impact of varying $\beta$ with on composed fused embeddings
Figure 3: Impact of changing $\alpha$/$\beta$ on Recall@5 performance across different PDV applications. For each row, results are shown for the CIReVL (left) and Pic2Word (right) baseline methods.
@@ 4.1 Effect of Using the PDV
Table 1: Average recall for different methods on Fashion-IQ validation dataset. † denotes that numbers are taken from the original paper.
@@
Table 2: Performance comparison on CIRCO and CIRR test datasets. As in previous works, for CIRCO, mAP@k is reported, while for CIRR both Recall@k and $R_{s}$@k metrics are used. † denotes that numbers are taken from the original paper.
@@
Analysing the PDV for Image (PDV-I) To evaluate whether $\Delta_{PDV}$ enhances the retrieval performance of image embeddings, we conducted experiments following the protocol described in Section 2. We modified image embeddings by adding $\Delta_{PDV}$ scaled with $\alpha$ values ranging from -0.5 to 2.0, where $\alpha=0$ represents the original image-only embeddings. As shown in Figure 3(b), Recall@K exhibits a positive correlation with $\alpha$ for values below 1. This upward trend continues until $\alpha=2.0$ for CIReVL, while Pic2Word’s performance peaks when $\alpha$ reaches 1.4. The performance of PDV-I was evaluated on the CIRR and CIRCO datasets by comparing it with other visual embedding-based methods, as detailed in Table 3. The results reveal that PDV-I achieved marginal improvements over existing approaches.
@@
Figure 4: Visualisation of the impact of $\alpha$/$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
@@ 4.2 ZS-CIR Benchmark Comparison
Dataset CIRCO CIRR
Metric mAP@k Recall@k $R_{s}$@k
Arch Method $\alpha_{I}$ k=5 k=10 k=25 k=50 k=1 k=5 k=10 k=50 k=1 k=2 k=3
Table 3: PDV-I performance on CIRCO and CIRR test datasets. Note that the image-only approach utilizes the visual embedding of the reference image, whereas the text-only approach employs the text embedding of the prompt.
We evaluated PDV-F alongside four baseline approaches (CIReVL, LDRE, Pic2Word, and SEARLE) across three benchmarks. Notably, CIReVL was tested with three different backbones on three datasets, as its models and intermediate results are publicly available. However, for the remaining methods, we conducted partial evaluations due to limited open-source availability or restricted support.
The numerical results are presented in Tables 1 and 2. On the FashionIQ benchmark, PDV-F yields substantial improvements for all baseline approaches, with CIReVL showing particularly strong gains that scale with backbone size. Similarly, all methods demonstrate significant performance improvements on CIRCO and CIRR datasets. Notably, CIReVL achieves larger improvements compared to other methods, with the most substantial gains observed when using small and medium backbone architectures. Our PDV-F implementation within the CIReVL framework consistently outperformed other state-of-the-art methods, including LinCIR and SEIZE, across most evaluation metrics. Similar to SEIZE, PDV-F offers the advantage of being entirely training-free; however, unlike SEIZE, it does not significantly increase feature extraction computational costs. While LinCIR demonstrates exceptional inference speed, it lacks the training-free nature of our approach, requiring dedicated model training before deployment.
5 Conclusion
@@
Appendix A Appendix
A.1 PDV Algorithm and Code

### v3

@@ Abstract
Zero-shot Composed Image Retrieval (ZS-CIR) enables image search using a reference image and a text prompt without requiring specialized text-image composition networks trained on large-scale paired data. However, current ZS-CIR approaches suffer from three critical limitations in their reliance on composed text embeddings: static query embedding representations, insufficient utilization of image embeddings, and suboptimal performance when fusing text and image embeddings. To address these challenges, we introduce the Prompt Directional Vector (PDV), a simple yet effective training-free enhancement that captures semantic modifications induced by user prompts. PDV enables three key improvements: (1) Dynamic composed text embeddings where prompt adjustments are controllable via a scaling factor, (2) composed image embeddings through semantic transfer from text prompts to image features, and (3) weighted fusion of composed text and image embeddings that enhances retrieval by balancing visual and semantic similarity. Our approach serves as a plug-and-play enhancement for existing ZS-CIR methods with minimal computational overhead. Extensive experiments across multiple benchmarks demonstrate that PDV consistently improves retrieval performance when integrated with state-of-the-art ZS-CIR approaches, particularly for methods that generate accurate compositional embeddings. The code will be released upon publication.
@@ 1 Introduction
Figure 1: Overview of Prompt Directional Vector (PDV) for Zero-Shot Composed Image Retrieval (ZS-CIR). (a) Standard ZS-CIR pipeline. (b) PDV calculation process. (c) Dynamic text embedding composition using PDV. (d) Fusion of composed text and image embeddings: PDV-modified composed image embedding combined with composed text embedding.
@@
Gap 1: Inefficient Iterative Search Due to Static Composed Embeddings Existing CIR studies primarily emphasize initial retrieval success but neglect users’ need to iteratively refine and steer results when initial searches fail. This limitation stems mainly from the static nature of composed text embeddings. Existing ZS-CIR approaches are unable to directly tune the composed text embedding, $\Psi_{T}(\mathcal{F}(I_{ref},P))$, to bring it closer to the target image embedding, $\Psi_{T}(\mathcal{F}(I_{ref},P))$. With existing ZS-CIR approaches users need to refine their prompts to iteratively improve retrieval results. However, this approach incurs additional manual effort for prompt construction and computational overhead from expensive feature extraction.
Gap 2: Underutilisation of Reference Image Embedding. Current methods generally do not utilize the embedding $\Psi_{I}(I_{ref})$ of the reference image directly for retrieval; instead, $\Psi_{I}(I_{ref})$ is used solely for composition. This omission stems from consistently poor retrieval performance (refer to Image-only results in Table S8) when incorporating these embeddings, as documented in multiple studies [22, 14, 3].
Gap 3: Suboptimal Performance of Image-Text Embedding Fusion. While the fusion of image and text embeddings outperforms single-modality approaches (image-only or text-only) [22, 14, 3], it still underperforms compared to composed text embeddings (refer to “Image+Text” results in Tables S5, S6 in the supplementary.).
@@
PDV: Addressing Gap 3. Lastly, several studies demonstrate that the direct fusion of image and text embeddings outperforms using either input feautre (image or text-prompt) alone [22, 14, 3]. However, this fusion approach still underperforms compared to using the composed text embeddings. This performance gap exists because prompt embeddings are significantly changed by incorporating context from the reference image. Specifically, $\Delta_{PDV}$ is not equivalent to $\Psi_{T}(P)$. To address this, we propose fusing the composed text and composed image embeddings, as illustrated in Figure 1d. Through varying the fusion weight factor $\beta$, we can dynamically control the balance between visual similarity to the reference image and semantic alignment with the prompt without needing to craft new prompts, or modify reference images. Lower $\beta$ values prioritize visual fidelity, while higher values emphasize semantic modifications specified in the prompt. The previous work, CompoDiff [12], also demonstrates similar controllability through adjusting visual and semantic weights. However, their method cannot be applied to other approaches as ours can, as our method is an add-on approach while theirs is specifically tailored to their own framework.
PDV serves as a plug-and-play enhancement for most ZS-CIR approaches, offering a simple and training-free solution. The computational overhead is minimal, requiring only the calculation of text and image embeddings from the reference image. We evaluate PDV by integrating it with four distinct ZS-CIR methods across various CIR benchmarks. Our experimental results demonstrate that all three use cases of PDV consistently improve upon baseline approaches, particularly when the baseline method already generates accurate compositional embeddings.
Contributions. Our main contributions are as follows: (i) We introduce the Prompt Directional Vector (PDV), a simple and training-free enhancement that overcomes limitations of current Zero-Shot CIR methods. (ii) We propose three novel applications of PDV: (1) dynamic composed text embedding synthesis through PDV scaling, which offers enhanced control over retrieval results without tedious prompt modification; (2) composed image embedding synthesis via semantic transfer of prompts to visual features through PDV addition, which prioritizes visual similarity; and (3) effective fusion of composed text and image embeddings, which improves overall performance and enables controllable balancing of visual and semantic similarity. (iii) Through extensive experiments on multiple benchmarks with four ZS-CIR methods, we demonstrate that PDV consistently improves retrieval performance with minimal computational overhead.
@@ 3.1 Baseline ZS-CIR Framework
If $I_{target}\notin\mathbb{I}_{top-k}$, the user must reformulate the prompt and repeat the feature extraction process to obtain alternative retrieval results, incurring time and computational resource costs. Notably, as shown in Eq. 1, only the composed feature embedding $\Psi_{T}(\mathcal{F}(I_{ref},P))$ directly influences the computation of $\mathbb{I}_{top-k}$. Although the gallery images are represented by their image embeddings, the image embedding of the reference image $\Psi_{I}(I_{ref})$ do not contribute to the retrieval process.
@@ 3.3 Strategies for Using PDV
When is PDV-T effective? In CIR, better retrieval performance is directly correlated with a smaller angle, $\theta$, between the target embedding vector $\Psi_{I}(I_{target})$ and the composed embedding vector $\Psi_{T}(\mathcal{F}(I_{ref},P))$ (refer to Figure 2(a)). When $\phi$, the angle between the calculated prompt directional vector $\Delta_{\mathrm{PDV}}$ and the ground truth prompt directional vector $\Delta_{\mathrm{GT}}$, is small, adjusting the parameter $\alpha$ can effectively reduce $\theta$. To demonstrate this relationship, we conducted a simulation based on a common real-world scenario where the magnitude of $\Delta_{\mathrm{PDV}}$ is smaller than $\Delta_{\mathrm{GT}}$ due to users writing less descriptive prompts (see Figure 2(a)). Based on simulation results (Figure 2(b)), when $\phi$ is less than 70 degrees, increasing $\alpha$ up to 3 reduces $\theta$. However, when $\phi$ is between 70 and 90 degrees, decreasing $\alpha$ is more effective in reducing $\theta$. These findings indicate that PDV’s performance is highly dependent on the accuracy of the baseline CIR method. With a strong baseline system, adjusting $\alpha$ becomes a straightforward and effective way to tune CIR results.
(a) Visualisation of $\theta$ and $\phi$.
(b) $\theta$ vs. $\alpha$ and $\phi$
Figure 2: A visualisation of how scaling $\alpha$ affects the $\theta$ angle between the composed embedding and the target embedding.
@@
Figure 3: Comparison of Image + Text (a) vs PDV-I (b).
3.4 Parameter Tuning and Efficient Searching
PDV introduces three hyperparameters: $\alpha_{T}$, $\alpha_{I}$, and $\beta$. Among these, $\alpha_{T}$ is most critical, and adjusting it alone typically suffices. Fine-tuning $\alpha_{I}$ and $\beta$ can further enhance retrieval performance and controllability. Detailed tuning guidelines are in Section S1.1111S represent supplementary material, and our automatic tuning method for $\alpha_{I}$ is described in Section S1.1.1.
PDV mainly reduces feature extraction costs in iterative search. We also find that filtering gallery images by prior ranking results lowers ranking costs, another key factor in CIR retrieval. Section S1.2 provides discussion and supporting experiments.
@@ 4.1 Effect of Using the PDV
Table 1: Average recall for different methods on Fashion-IQ validation dataset. Yellow highlighting shows PDV improvements over baseline. Orange = PDV improvements, bold = best, underline = second. †Numbers from original paper.
@@
Table 2: Performance comparison on CIRCO and CIRR test datasets. As in previous works, for CIRCO, mAP@k is reported, while for CIRR both Recall@k and $R_{s}$@k metrics are used. Orange = PDV improvements, bold = best, underline = second. †Numbers from original paper.
@@
(a) PDV-T: Impact of $\alpha$ scaling on composed text embeddings
(b) PDV-I: Impact of $\alpha$ scaling on composed image embeddings
(c) PDV-F: Impact of varying $\beta$ with on composed fused embeddings
Figure 4: Impact of changing $\alpha$$\beta$ on Recall@5 performance across different PDV applications. For each row, results are shown for the CIReVL (left) and Pic2Word (right) baseline methods.
@@
Analysing the PDV for Image (PDV-I) To evaluate whether $\Delta_{PDV}$ enhances the retrieval performance of image embeddings, we conducted experiments following the protocol described in Section 2. We modified image embeddings by adding $\Delta_{PDV}$ scaled with $\alpha$ values ranging from -0.5 to 2.0, where $\alpha=0$ represents the original image-only embeddings. As shown in Figure 4(b), Recall@K exhibits a positive correlation with $\alpha$ for values below 1. This upward trend continues until $\alpha=2.0$ for CIReVL, while Pic2Word’s performance peaks when $\alpha$ reaches 1.4. The performance of PDV-I was evaluated on the FashionIQ, CIRR and CIRCO datasets by comparing it with other visual embedding-based methods, as detailed in Tables S5, S6 in the supplementary material. The results reveal that PDV-I achieved significant improvements over existing approaches.
@@
Figure 5: Visualisation of the impact of $\alpha$$\beta$ scaling on top-5 retrieval results. CIReVL with ViT-B-32 Clip model is the baseline method used. Representative examples with prompts from three datasets: FashionIQ (left), CIRR (middle), and CIRCO (right) are shown at the top. Green and blue bounding boxes indicate true positives and near-true positives, respectively.
@@ 4.2 ZS-CIR Benchmark Comparison
We evaluated PDV-F alongside four baseline approaches (CIReVL, LDRE, Pic2Word, and SEARLE) across three benchmarks. Notably, CIReVL was tested with three different backbones on three datasets, as its models and intermediate results are publicly available. However, for the remaining methods, we conducted evaluations using the models which are publicly available.
Numerical results are presented in Tables 1 and 2. The hyperparameters shown in these tables were selected based on the valid ranges identified in Figure 4. Since these parameters are intended for user adjustment and users are expected to tune them in response to initial retrieval results, we present the optimal hyperparameters from the three trials. On the FashionIQ benchmark, PDV-F yields substantial improvements for all baseline approaches, with CIReVL showing particularly strong gains that scale with backbone size. Similarly, all methods demonstrate significant performance improvements on CIRCO and CIRR datasets. Notably, CIReVL achieves larger improvements compared to other methods, with the most substantial gains observed when using small and medium backbone architectures. Our PDV-F implementation within the CIReVL framework consistently outperformed other state-of-the-art methods, including LinCIR and SEIZE, across most evaluation metrics. Similar to SEIZE, PDV-F offers the advantage of being entirely training-free; however, unlike SEIZE, it does not significantly increase feature extraction computational costs. While LinCIR demonstrates exceptional inference speed, it lacks the training-free nature of our approach, requiring dedicated model training before deployment.
5 Conclusion & Future Applications
@@
Supplement A Additional Experiments and Results
A.1 Regarding Hyper-parameter Tuning
In this work, we introduce three parameters: $\alpha_{I}$, $\alpha_{T}$, and $\beta$. Once their semantic roles are understood, manually adjusting them becomes intuitive and straightforward:
$\alpha_{T}\rightarrow$ Influence of the text prompt on the semantic content of the reference image
$\alpha_{I}\rightarrow$ Influence of the text prompt on the visual content of the reference image
$\beta\rightarrow$ Controls trade-off between visual and semantic content
We also examined the potential for automatic tuning. However, due to several uncertain factors—such as prompt quality, baseline performance on the target dataset, and user expectations—it remains highly challenging to optimise all three parameters automatically. Nevertheless, we find that $\alpha_{I}$ can be tuned automatically to reduce the gap between PDV-I and PDV-T. In the following subsections, we discuss the manual adjustment of each parameter and the automatic tuning of $\alpha_{I}$.
A.1.1 Tuning $\alpha_{T}$: Influence of Text Prompt
Among these parameters, the most important is $\alpha_{T}$, which primarily controls the influence of the text prompt. If a user observes that the semantic changes in the top retrieved results are insufficient, $\alpha_{T}$ should be increased; conversely, if the changes are too strong, it should be decreased. For example, in the top case of Figure S6, when $\alpha_{T}=1$, the skirt does not yet display clear white stripes. Increasing $\alpha_{T}$ produces results with more distinct white stripes. In contrast, in the middle and bottom examples of Figure S6, the retrieved results with $\alpha_{T}=1$ are already valid. Further increasing $\alpha_{T}$ in these cases makes the semantic changes too strong, leading the method to return invalid results.
A.1.2 Manual Tuning $\alpha_{I}$: Influence of Text Prompt
The role of $\alpha_{I}$ is similar to that of $\alpha_{T}$. It also controls the strength of the prompt, but in this case, the composition is with the original visual embedding $\Psi_{I}(I_{ref})$. When $\alpha_{I}=0$, PDV-I reduces to content-based image retrieval. From the ablation results shown in Figure S13, we observe that setting $\alpha_{I}=1$ is generally safe, as most methods achieve consistent improvements when $\alpha_{I}$ is increased from 0.5-0.5. Nevertheless, further increases in $\alpha_{I}$ can also be beneficial. Users should continue increasing $\alpha_{I}$ when the top retrieved results are overly similar to the reference image and fail to incorporate the semantic concepts specified in the prompt. For instance, in the top and bottom examples of Figure S7, when $\alpha_{I}>1$, the top-1 retrieval results successfully present the semantic elements described in the user prompt.
A.1.3 Tuning $\beta$: Fusion Factor
The parameter $\beta$ is used in PDV-F, which fuses PDV-I and PDV-T. Its value ranges from 0 to 1. When $\beta=1$, PDV-F is equivalent to PDV-T, and when $\beta=0$, it reduces to PDV-I. From the ablation results shown in Figure S14, we observe that most methods achieve improved performance when $\beta$ lies between 0.6 and 0.9. Beyond performance optimization, $\beta$ also plays a crucial role in balancing retrieval characteristics. As illustrated in Figure S8, lower $\beta$ values ($\beta<0.5$) emphasize visual similarity, producing top results that closely resemble the reference image $I_{ref}$ in terms of appearance. In contrast, higher $\beta$ values ($\beta>0.5$) prioritize semantic alignment, incorporating conceptual elements described in the text prompt. This provides fine-grained control over whether the retrieval system favors visual fidelity or semantic relevance.
Figure S6: Qualitative results of PDV-T showing the effect of different $\alpha_{T}$ values. For each query, we display the top-1 retrieval result for three different $\alpha_{T}$ settings. The middle result uses $\alpha_{T}=1$ (baseline), the left result uses a smaller $\alpha_{T}$ value, and the right result uses a larger $\alpha_{T}$ value. All $\alpha_{T}$ values are within the range [0.5,2][-0.5,2].
Figure S7: Qualitative results of PDV-I showing the effect of different $\alpha_{I}$ values. For each query, we display the top-1 retrieval result for three different $\alpha_{I}$ settings. The middle result uses $\alpha_{I}=1$ (baseline), the left result uses a smaller $\alpha_{I}$ value, and the right result uses a larger $\alpha_{I}$ value. All $\alpha_{I}$ values are within the range [0.5,2][-0.5,2].
Figure S8: Qualitative results of PDV-F illustrating the effect of different $\beta$ values. For each query, we show the top-1 retrieval result under three settings: $\beta=0$ (left), $\beta=0.5$ (middle), and $\beta=1$ (right).
A.1.4 Automatically Tuning $\alpha_{I}$
Based on the experimental results, we observe that PDV-T consistently outperforms PDV-I. If the features of PDV-I, denoted as $\Phi_{\text{PDV-I}}$, are more closely aligned with those of PDV-T, $\Phi_{\text{PDV-T}}$, the performance of PDV-I can approach that of PDV-T. Motivated by this observation, we tune the parameter $\alpha_{I}$ (while keeping $\Phi_{\text{PDV-T}}$ fixed) to minimize the $\ell_{2}$ distance between $\Phi_{\text{PDV-I}}$ and $\Phi_{\text{PDV-T}}$, as expressed in Equation S5. To determine the optimal value of $\alpha_{I}$, we employ the Nelder–Mead optimization method [19], which is a derivative-free and straightforward approach, making it particularly convenient to implement.
$\alpha_{I}=\underset{\alpha}{\arg\min}\;\mathcal{L}\!\left(\Phi_{\text{PDV-T}},\,\Phi_{\text{PDV-I}}(\alpha)\right),$ (S5)
To evaluate the effectiveness of the proposed method, we fix $\alpha_{T}=1$, making PDV-T equivalent to the baseline, and compare the performance of tuned $\alpha_{I}$ against the fixed setting $\alpha_{I}=1$. We conduct experiments on the FashionIQ dataset using three different methods with multiple backbone architectures. As shown in Table S3, our approach successfully determines customized $\alpha_{I}$ values for each setting. In all cases, R@50 shows consistent improvements over the baseline, with the largest gain of 23% achieved by CIReVL with the ViT-B/32 backbone on the Toptee subset. R@10 also improves steadily in most scenarios, particularly with the CIReVL method. However, for Pic2Word, R@10 decreases by 3.42% on the Dress subset.
Backbone Method $\alpha_{I}$ Shirt Dress Toptee
R@10 R@50 R@10 R@50 R@10 R@50
ViT-B/32 SEARLE 1.57/1.65/1.58 15.68% 15.62% 20.04% 20.51% 14.52% 11.18%
CIReVL 1.92/2.24/2.02 25.65% 18.82% 24.95% 16.86% 27.81% 23.11%
ViT-L/14 CIReVL 1.90/2.16/1.95 12.96% 9.58% 15.12% 10.50% 17.70% 12.80%
Pic2Word 1.47/1.46/1.48 0.00% 3.09% -3.42% 2.07% 0.00% 0.74%
SEARLE 1.67/1.82/1.73 5.84% 10.62% 16.57% 9.41% 6.15% 9.86%
ViT-G/14 CIReVL 1.50/1.63/1.53 10.43% 6.61% 19.15% 14.30% 17.51% 12.14%
Table S3: Performance differences with automatic $\alpha_{I}$ tuning compared to the fixed setting $\alpha_{I}=1$ on the FashionIQ datasets. The $\alpha_{I}$ column reports the tuned values for the Shirt, Dress, and Toptee subsets.
A.2 Efficient Retrieval with PDV
PDV is designed to enhance the retrieval performance of baseline methods in a subsequent search, triggered when an initial query fails, without incurring the high computational cost typically associated with iterative search processes.
The computational bottleneck in Zero-Shot Composed Image Retrieval (ZS-CIR) systems stems from two primary operations: feature extraction and similarity ranking.
Regarding feature extraction, the cost is dictated by the model employed. Recent ZS-CIR approaches rely on large vision-language models, whose feature extraction overhead is significant, as detailed in Table S4. In contrast, PDV generates new features for subsequent trials by building upon the embeddings from the initial retrieval. This process involves only efficient scalar multiplications and matrix additions, making its per-trial feature extraction cost nearly negligible. The only substantial computational overhead is the one-time initial calculation of the reference image embedding, $\Psi_{I}(I_{ref})$.
The cost of similarity ranking, on the other hand, is primarily a function of the feature dimension and the gallery size. While the feature dimension is fixed by the base model, the gallery size can be reduced for subsequent searches. To improve efficiency, we integrate a simple filtering strategy with PDV: items whose distance from the query exceeds a predefined threshold are removed from the gallery for subsequent ranking. While not unique to PDV, we believe this is the first discussion of such an optimization in a ZS-CIR context.
We evaluated this approach on the FashionIQ dataset using two baseline methods, CIReVL and Pic2Word. As shown in Table S5, for CIReVL, a threshold of 0.8 filters out over 80% of the gallery items while degrading the R@50 metric by at most 1.31%. For Pic2Word, a threshold of 0.75 filters out over 68% of the gallery with a maximum R@50 decrease of 3.73%. These results demonstrate that PDV, combined with gallery filtering, offers a highly effective trade-off, significantly accelerating retrieval speed while maintaining competitive accuracy.
Method Feature Extraction Time (Sec.)
Initial Retrial
Pic2Word 0.02 0.02
+ PDV 0.03 0.00
LinCIR 0.02 0.02
+ PDV 0.03 0.00
KEDs 0.03 0.04
CIReVL (2 Captions) 1.23 1.23
+ PDV 1.24 0.00
LDRE (20 Captions) 17.30 17.30
+ PDV 17.31 0.00
Table S4: Comparison of computation efficiency of the baseline ZS-CIR approaches on NVIDIA A100 GPU.
Backbone Method Dataset Threshold Filtered Ratio Change in R@10 Change in R@50
ViT-B/32 CIReVL + PDV-F Toptee 0.8 90.85% -2.88% -1.31%
Dress 0.8 82.31% 0.00% -0.73%
Shirt 0.8 87.96% -1.22% -1.31%
ViT-L/14 Pic2Word + PDV-F Toptee 0.75 85.62% -1.24% -0.94%
Dress 0.75 68.79% 0.00% 0.12%
Shirt 0.75 88.36% -1.91% -3.73%
Table S5: Performance changes of PDV-F after applying filtering to the initial retrieval results on the FashionIQ dataset.
A.3 $\phi$ Angles of the Baselines
In Section 3.3 of the main paper, we discussed that PDV’s performance is highly correlated with baseline performance and provided theoretical justification through simulation results as shown in Figure 2. We have introduced a new parameter $\phi$, which is the angle between the calculated prompt directional vector $\Delta_{\mathrm{PDV}}$ and the ground truth prompt directional vector $\Delta_{\mathrm{GT}}$. When $\phi$ is small, adjusting the parameter $\alpha$ can effectively reduce $\theta$ which is the angle between the target embedding vector $\Psi_{I}(I_{target})$ and the composed embedding vector $\Psi_{T}(\mathcal{F}(I_{ref},P))$.
Here, we present the actual $\phi$ values for three baseline methods—CIReVL, Pic2Word, and SEARLE—using different backbones across three subdatasets of the FashionIQ dataset. Figure S9 presents the $\phi$ values for three baseline methods—CIReVL, Pic2Word, and SEARLE—across the FashionIQ subdatasets. While we observed the expected trend of stronger models exhibiting smaller $\phi$ values, we were surprised to find that the state-of-the-art CIReVL baseline maintains a large $\phi$ angle of approximately $65^{\circ}$. This finding suggests that PDV has not yet been evaluated with optimal baseline models, despite already yielding considerable gains with CIReVL. We therefore anticipate that PDV will be even more effective when paired with future baselines that achieve a smaller $\phi$ (ideally $60^{\circ}$), the regime where, according to Figure 2(b), our method is most impactful.
Figure S9: Performance Comparison: $\phi$ Angles by Method and Model Across FashionIQ Datasets
A.4 Additional Quantitative Results
In this supplementary section, we present additional quantitative results that were omitted from the main paper due to space constraints.
A.4.1 Ablation Analysis
While Figure 3 in the main paper illustrates the effects of scaling factor $\alpha$ and fusion factor $\beta$ on Recall@5 performance across various PDV applications, Figures S12, S13, and S14 present complementary results for Recall@10 and Recall@50 metrics.
A.4.2 PDV-I Results
We also provide additional PDV-I results achieved on the validation set of the FashionIQ dataset, as shown in Tables S7 and S8. PDV-I also achieved significant improvements over existing approaches that directly leverage image embeddings for retrieval.
Lastly, we provide a detailed visualization of the impact of $\alpha$$\beta$ scaling on top-5 retrieval results. Figure S11 illustrates the performance of CIReVL with the ViT-B-32 CLIP model across three different datasets.
A.4.3 Zero-shot methods with PDV Vs. Supervised Methods
We compared state-of-the-art zero-shot composed image retrieval (ZS-CIR) methods enhanced with PDV against supervised methods on the FashionIQ and CIRR datasets. Our evaluation included early supervised methods such as TIRG [26] and ARTEMIS citedelmas2022artemis, as well as recent state-of-the-art approaches like CCIN [24] and SPRC [2]. The comparative results presented in Table S6 demonstrate that PDV achieves remarkably competitive performance against supervised methods. The PDV-based approaches significantly outperform early supervised baselines, with substantial improvements over TIRG (41.90% vs 14.13% R@10 on FashionIQ Dress) and ARTEMIS [9] (41.90% vs 25.68%). While PDV methods do not quite match the performance of recent state-of-the-art approaches like CCIN and SPRC, the performance gap remains relatively modest—typically within 7-9 percentage points on FashionIQ and approximately 8-9 points on CIRR’s mean score (72.85% vs 81.66% for CCIN). This narrow performance gap is particularly noteworthy given that ZS-CIR methods with PDV operate without human-annotated training data. These results suggest that unsupervised approaches are reaching a level of effectiveness that positions them as viable alternatives to supervised methods.
Method FashionIQ CIRR
Dress Shirt Toptee Mean
R@10 R@50 R@10 R@50 R@10 R@50 R@1 R@5 R@10 R@50 Mean
TIRG [26] 14.13 34.61 13.10 30.91 14.79 34.37 23.66 14.61 48.37 64.08 90.03 54.27
ARTEMIS [9] 25.68 51.05 21.57 44.13 25.89 55.06 37.68 16.96 46.10 61.31 87.73 53.03
CLIP4CIR [5] 33.81 59.40 39.99 60.45 41.41 65.37 50.03 38.53 69.98 81.86 95.93 71.58
CompoDiff [12] 40.65 57.14 36.87 57.39 43.93 61.17 49.53 22.35 54.36 73.41 91.77 60.47
TG-CIR [27] 45.22 69.66 52.60 72.52 56.14 77.10 58.05 45.25 78.29 87.16 97.30 77.00
SPRC [2] 48.83 72.09 53.83 74.14 58.13 78.58 64.27 51.96 82.12 89.74 97.69 80.37
CCIN [24] 49.38 72.58 55.93 74.14 57.93 77.56 64.59 53.41 84.05 91.17 98.00 81.66
CIReVL + PDV 41.90 58.19 40.70 62.82 48.09 67.77 53.25 38.15 67.93 77.90 92.77 69.19
LDRE + PDV - - - - - - - 42.51 72.22 81.71 94.94 72.85
Table S6: Comparison of PDV with supervised methods on FashionIQ and CIRR datasets.
Supplement B PDV Algorithm and Code